OpenStudy (anonymous):

Solve log11 (y + 8) + log11 4 = log11 60

OpenStudy (anonymous):

log11 (y + 8) + log11 4 = log11 60 am assuming 11 is a base, so log11 (y + 8) = log11 60- log11 4 remember when you subtract numbers you divide logs: log11 (y + 8) = log11 (60/4) log11 (y + 8) = log11 15 now just ignore the log11 y + 8 = 15 y = 7

hero (hero):

You can also do it this way: \(\log_{11} {4(y+8)} = \log_{11}{60}\) then set 4(y+8) = 60 and solve for y